C=500+0.4Yd, MPS = 0.6, I = 500, G = 700, X = 200 and M=50+0.2Y the Tax rate is 20% of Y. The population is 90 people. Calculate NI equilibrium *
1
Expert's answer
2021-09-22T07:54:50-0400

"Y=C+I+G+X-M\\Y=500+0.4(Y-0.2)+500+700+200-50+0.6Y\\Y=1849.92+0.6Y\\0.4Y=1849.92\\Y=\frac{1849.92}{0.4}\\Y=4624.8"


Therefore the national income equilibrium is 4624.80